# A note on division rings satisfying generalized rational identities with anti-automorphisms

### v2

@@ 3. Proof of the main result
Proof.
For an original proof of the equivalence between (1) and (2), we refer to [11, Theorem 1]. In addition, it is clear that every PI is, in particular, a GPI, and every GPI is a GRI, so we have the implications from (2) to (3) and from (3) to (4). A proof that (4) implies (1) can be found in [6]. Therefore, assertions (1)–(4) are all equivalent. ∎
@@
Proof.  According to Corollary 2.3, if $D$ satisfies a $\sigma^{m}$-GPI, then we can assume that $D$ satisfies a $\sigma^{m}$-linear polynomial identity
$f(x_{1},x_{2},\cdots,x_{n},x_{1}^{\sigma},x_{2}^{\sigma},\cdots,x_{n}^{\sigma},\cdots,x_{1}^{\sigma^{m}},x_{2}^{\sigma^{m}},\cdots,x_{n}^{\sigma^{m}})\in D\left\langle X_{m}\right\rangle.$
Then, for all $r_{1},r_{2},\cdots,r_{n}\in D$, we have
$f(r_{1},r_{2},\cdots,r_{n},r_{1}^{\sigma},r_{2}^{\sigma},\cdots,r_{n}^{\sigma},\cdots,r_{1}^{\sigma^{m}},r_{2}^{\sigma^{m}},\cdots,r_{n}^{\sigma^{m}})=0.$
@@
$f_{1}(x_{1},x_{1}^{\sigma},\cdots,x_{1}^{\sigma^{m}})=f(x_{1},r_{2},\cdots,r_{n},x_{1}^{\sigma},r_{2}^{\sigma},\cdots,r_{n}^{\sigma},\cdots,x_{1}^{\sigma^{m}},r_{2}^{\sigma^{m}},\cdots,r_{n}^{\sigma^{m}}).$
If $f_{1}(x_{1},x_{1}^{\sigma},\cdots,x_{1}^{\sigma^{m}})\neq 0$, then by using Lemma 2.4, $\dim_{F}D<\infty$. Therefore, according to Lemma 3.1, the division ring $D$ satisfies a GPI. Suppose $f_{1}(x_{1},x_{1}^{\sigma},\cdots,x_{1}^{\sigma^{m}})=0$. For all $r_{1},r_{2},\cdots,r_{n},r_{11},r_{12},\cdots,r_{1m}\in D$, we have
$f(r_{1},r_{2},\cdots,r_{n},r_{11},r_{2}^{\sigma},\cdots,r_{n}^{\sigma},\cdots,r_{1m},r_{2}^{\sigma^{m}},\cdots,r_{n}^{\sigma^{m}})=0.$
@@
$f_{2}(x_{2},x_{2}^{\sigma},\cdots,x_{2}^{\sigma^{m}})$
$f(r_{1},x_{2},r_{3}\cdots,r_{n},r_{11},x_{2}^{\sigma},r_{3}^{\sigma},\cdots,r_{n}^{\sigma},\cdots,r_{1m},x_{2}^{\sigma^{m}},r_{3}^{\sigma^{m}},\cdots,r_{n}^{\sigma^{m}}).$
If $f_{2}(x_{2},x_{2}^{\sigma},\cdots,x_{2}^{\sigma^{m}})\neq 0$, then $D$ satisfies a GPI. Conversely, if
$f_{2}(x_{2},x_{2}^{\sigma},\cdots,x_{2}^{\sigma^{m}})=0$
then for all $r_{1},r_{2},\cdots,r_{n},r_{11},r_{12},\cdots,r_{1m},r_{21},r_{22},\cdots,r_{2m}\in D$, we have
$f(r_{1},r_{2},\cdots,r_{n},r_{11},r_{21},\cdots,r_{n}^{\sigma},\cdots,r_{1m},r_{2m},\cdots,r_{n}^{\sigma^{m}})=0.$
By repeating the same method for $x_{3},x_{4},\cdots,x_{n}$, we can assume that for all $r_{1},r_{2},\cdots,r_{nm}\in D$,
$f(r_{1},r_{2},\cdots,r_{11},r_{21},\cdots,r_{n1},\cdots,r_{1m},r_{2m},\cdots,r_{nm})=0.$
@@
Let $F$ be a field and $R$ be a ring whose center contains $F$. If $f(t)=a_{n}t^{n}+a_{n-1}t^{n-1}+\dots+a_{1}t+a_{0}$ is a nonzero polynomial over $R$, then $f(t)=0$ has at most $n$ solutions in $F$.
Proof.
This lemma is a corollary of [16, Proposition 2.3.27]. ∎
@@
$f=f(x_{1},\cdots,x_{n},x_{1}^{\sigma},\cdots,x_{n}^{\sigma},\cdots,x_{1}^{\sigma^{m}},\cdots,x_{n}^{\sigma^{m}})$
@@
$(r_{1},\cdots,r_{n},r_{1}^{\sigma},\cdots,r_{n}^{\sigma},\cdots,r_{1}^{\sigma^{m}},\cdots,r_{n}^{\sigma^{m}}).$
@@
$(r_{1}+x_{1}t,\cdots,r_{n}+x_{n}t,r_{1}^{\sigma}+x_{1}^{\sigma}t,\cdots,r_{n}^{\sigma}+x_{n}^{\sigma}t,\cdots,r_{1}^{\sigma^{m}}+x_{1}^{\sigma^{m}}t,\cdots,r_{n}^{\sigma^{m}}+x_{n}^{\sigma^{m}}t)$
@@
$f(r_{1}+x_{1}t,\cdots,r_{n}+x_{n}t,r_{1}^{\sigma}+x_{1}^{\sigma}t,\cdots,r_{n}^{\sigma}+x_{n}^{\sigma}t,\cdots,r_{1}^{\sigma^{m}}+x_{1}^{\sigma^{m}}t,\cdots,r_{n}^{\sigma^{m}}+x_{n}^{\sigma^{m}}t)$
$=f_{0}+\sum_{i=1}^{\infty}f_{i}(x_{1},\cdots,x_{n},x_{1}^{\sigma},\cdots,x_{n}^{\sigma},\cdots,x_{1}^{\sigma^{m}},\cdots,x_{n}^{\sigma^{m}})t^{i}$
where $f_{0}=f(r_{1},\cdots,r_{n},r_{1}^{\sigma},\cdots,r_{n}^{\sigma},\cdots,r_{1}^{\sigma^{m}},\cdots,r_{n}^{\sigma^{m}})$ and $f_{i}$ $(i\geq 1)$ are homogeneous polynomials of degree $i$ in $D\langle X_{m}\rangle$. Moreover, there exists $i_{0}$ such that $f_{i_{0}}\neq 0$. We have
$f(r_{1}+d_{1}t,\cdots,r_{n}+d_{n}t,r_{1}^{\sigma}+d_{1}^{\sigma}t,\cdots,r_{n}^{\sigma}+d_{n}^{\sigma}t,\cdots,r_{1}^{\sigma^{m}}+d_{1}^{\sigma^{m}}t,\cdots,r_{n}^{\sigma^{m}}+d_{n}^{\sigma^{m}}t)$
$=\sum_{i=1}^{\infty}f_{i}(d_{1},\cdots,d_{n},d_{1}^{\sigma},\cdots,d_{n}^{\sigma},\cdots,d_{1}^{\sigma^{m}},\cdots,d_{n}^{\sigma^{m}})t^{i}.$
Since $f(r_{1}+d_{1}t,\cdots,r_{n}+d_{n}t,r_{1}^{\sigma}+d_{1}^{\sigma}t,\cdots,r_{n}^{\sigma}+d_{n}^{\sigma}t,\cdots,r_{1}^{\sigma^{m}}+d_{1}^{\sigma^{m}}t,\cdots,r_{n}^{\sigma^{m}}+d_{n}^{\sigma^{m}}t)$ is an element in the quotient ring $D(t)$ of the polynomial ring $D[t]$, it can be represented as
$f(r_{1}+d_{1}t,\cdots,r_{n}+d_{n}t,r_{1}^{\sigma}+d_{1}^{\sigma}t,\cdots,r_{n}^{\sigma}+d_{n}^{\sigma}t,\\ \cdots,r_{1}^{\sigma^{m}}+d_{1}^{\sigma^{m}}t,\cdots,r_{n}^{\sigma^{m}}+d_{n}^{\sigma^{m}}t)=\frac{g(t)}{h(t)}$
where $g(t),h(t)\in D[t]$ and $h(t)\neq 0$. Since the field $F$ is infinite, there exist infinitely many $p\in F$ such that $f$ is defined at $(r_{1}+d_{1}p,\cdots,r_{n}+d_{n}p)$. Furthermore, since $f$ is a $\sigma^{m}$-GRI of $D$, we have
$f(r_{1}+d_{1}p,\cdots,r_{n}+d_{n}p,r_{1}^{\sigma}+d_{1}^{\sigma}p,\cdots,r_{n}^{\sigma}+d_{n}^{\sigma}p,\cdots,r_{1}^{\sigma^{m}}+d_{1}^{\sigma^{m}}p,\cdots,r_{n}^{\sigma^{m}}+d_{n}^{\sigma^{m}}p)$
$=\frac{g(p)}{h(p)}=0,$
@@
$f(r_{1}+d_{1}t,\cdots,r_{n}+d_{n}t,r_{1}^{\sigma}+d_{1}^{\sigma}t,\cdots,r_{n}^{\sigma}+d_{n}^{\sigma}t,\cdots,r_{1}^{\sigma^{m}}+d_{1}^{\sigma^{m}}t,\cdots,r_{n}^{\sigma^{m}}+d_{n}^{\sigma^{m}}t)=0,$
@@
$=\sum_{i=1}^{\infty}f_{i}(d_{1},\cdots,d_{n},d_{1}^{\sigma},\cdots,d_{n}^{\sigma},\cdots,d_{1}^{\sigma^{m}},\cdots,d_{n}^{\sigma^{m}})t^{i}=0.$
Therefore, $f_{i}(d_{1},d_{2},\cdots,d_{n},d_{1}^{\sigma},d_{2}^{\sigma},\cdots,d_{n}^{\sigma},\cdots,d_{1}^{\sigma^{m}},d_{2}^{\sigma^{m}},\cdots,d_{n}^{\sigma^{m}})=0$ for all $i$. Consequently, for all $d_{1},d_{2},\cdots,d_{n}\in D$, we have
$f_{i_{0}}(d_{1},d_{2},\cdots,d_{n},d_{1}^{\sigma},d_{2}^{\sigma},\cdots,d_{n}^{\sigma},\cdots,d_{1}^{\sigma^{m}},d_{2}^{\sigma^{m}},\cdots,d_{n}^{\sigma^{m}})=0.$
@@
$f_{i_{0}}(x_{1},x_{2},\cdots,x_{n},x_{1}^{\sigma},x_{2}^{\sigma},\cdots,x_{n}^{\sigma},\cdots,x_{1}^{\sigma^{m}},x_{2}^{\sigma^{m}},\cdots,x_{n}^{\sigma^{m}})$
is a $\sigma^{m}$-GPIGPI of $D$. According to Proposition 3.2, the ring $D$ satisfies a GPI. Therefore, $D$ is centrally finite by Lemma 3.1. ∎

### v3

@@ 3 Proof of the main result
Proof 3.2.
For an original proof of the equivalence between (1) and (2), we refer to [11, Theorem 1]. In addition, it is clear that every PI is, in particular, a GPI, and every GPI is a GRI, so we have the implications from (2) to (3) and from (3) to (4). A proof that (4) implies (1) can be found in [6]. Therefore, assertions (1)–(4) are all equivalent.
@@
Proof 3.4.
According to Corollary 2.4, if $D$ satisfies a $\sigma^{m}$-GPI, then we can assume that $D$ satisfies a $\sigma^{m}$-linear polynomial identity
$f(x_{1},x_{2},\dots,x_{n},x_{1}^{\sigma},x_{2}^{\sigma},\dots,x_{n}^{\sigma},\dots,x_{1}^{\sigma^{m}},x_{2}^{\sigma^{m}},\dots,x_{n}^{\sigma^{m}})\in D\left\langle X_{m}\right\rangle.$
Then, for all $r_{1},r_{2},\dots,r_{n}\in D$, we have
$f(r_{1},r_{2},\dots,r_{n},r_{1}^{\sigma},r_{2}^{\sigma},\dots,r_{n}^{\sigma},\dots,r_{1}^{\sigma^{m}},r_{2}^{\sigma^{m}},\dots,r_{n}^{\sigma^{m}})=0.$
@@
$f_{1}(x_{1},x_{1}^{\sigma},\dots,x_{1}^{\sigma^{m}})=f(x_{1},r_{2},\dots,r_{n},x_{1}^{\sigma},r_{2}^{\sigma},\dots,r_{n}^{\sigma},\dots,x_{1}^{\sigma^{m}},r_{2}^{\sigma^{m}},\dots,r_{n}^{\sigma^{m}}).$
If $f_{1}(x_{1},x_{1}^{\sigma},\dots,x_{1}^{\sigma^{m}})\neq 0$, then by using Lemma 2.5, $\dim_{F}D<\infty$. Therefore, according to Lemma 3.1, the division ring $D$ satisfies a GPI. Suppose $f_{1}(x_{1},x_{1}^{\sigma},\dots,x_{1}^{\sigma^{m}})=0$. For all $r_{1},r_{2},\dots,r_{n},r_{11},r_{12},\dots,r_{1m}\in D$, we have
$f(r_{1},r_{2},\dots,r_{n},r_{11},r_{2}^{\sigma},\dots,r_{n}^{\sigma},\dots,r_{1m},r_{2}^{\sigma^{m}},\dots,r_{n}^{\sigma^{m}})=0.$
@@
$f_{2}(x_{2},x_{2}^{\sigma},\dots,x_{2}^{\sigma^{m}})=f(r_{1},x_{2},r_{3}\dots,r_{n},r_{11},x_{2}^{\sigma},r_{3}^{\sigma},\dots,r_{n}^{\sigma},\dots,r_{1m},x_{2}^{\sigma^{m}},r_{3}^{\sigma^{m}},\dots,r_{n}^{\sigma^{m}}).$
If $f_{2}(x_{2},x_{2}^{\sigma},\dots,x_{2}^{\sigma^{m}})\neq 0$, then $D$ satisfies a GPI. Conversely, if
$f_{2}(x_{2},x_{2}^{\sigma},\dots,x_{2}^{\sigma^{m}})=0$
then for all $r_{1},r_{2},\dots,r_{n},r_{11},r_{12},\dots,r_{1m},r_{21},r_{22},\dots,r_{2m}\in D$, we have
$f(r_{1},r_{2},\dots,r_{n},r_{11},r_{21},\dots,r_{n}^{\sigma},\dots,r_{1m},r_{2m},\dots,r_{n}^{\sigma^{m}})=0.$
By repeating the same method for the variables $x_{3},x_{4},\dots,x_{n}$, we can assume that for all $r_{1},r_{2},\dots,r_{nm}\in D$,
$f(r_{1},r_{2},\dots,r_{11},r_{21},\dots,r_{n1},\dots,r_{1m},r_{2m},\dots,r_{nm})=0.$
@@
Let $F$ be a field and $R$ be a ring whose center contains $F$. If
$f(t)=a_{n}t^{n}+a_{n-1}t^{n-1}+\dots+a_{1}t+a_{0}$
is a nonzero polynomial over $R$, then $f(t)=0$ has at most $n$ solutions in $F$.
Proof 3.6.
This lemma is a corollary of [16, Proposition 2.3.27].
@@
$f=f(x_{1},\dots,x_{n},x_{1}^{\sigma},\dots,x_{n}^{\sigma},\dots,x_{1}^{\sigma^{m}},\dots,x_{n}^{\sigma^{m}})$
@@
$(r_{1},\dots,r_{n},r_{1}^{\sigma},\dots,r_{n}^{\sigma},\dots,r_{1}^{\sigma^{m}},\dots,r_{n}^{\sigma^{m}}).$
@@
$(r_{1}+x_{1}t,\dots,r_{n}+x_{n}t,r_{1}^{\sigma}+x_{1}^{\sigma}t,\dots,r_{n}^{\sigma}+x_{n}^{\sigma}t,\dots,r_{1}^{\sigma^{m}}+x_{1}^{\sigma^{m}}t,\dots,r_{n}^{\sigma^{m}}+x_{n}^{\sigma^{m}}t)$
@@
$\ f(r_{1}+x_{1}t,\dots,r_{n}+x_{n}t,r_{1}^{\sigma}+x_{1}^{\sigma}t,\dots,r_{n}^{\sigma}+x_{n}^{\sigma}t,\dots,r_{1}^{\sigma^{m}}+x_{1}^{\sigma^{m}}t,\dots,r_{n}^{\sigma^{m}}+x_{n}^{\sigma^{m}}t)$
$=$ $\ f_{0}+\sum_{i=1}^{\infty}f_{i}(x_{1},\dots,x_{n},x_{1}^{\sigma},\dots,x_{n}^{\sigma},\dots,x_{1}^{\sigma^{m}},\dots,x_{n}^{\sigma^{m}})t^{i}$
where $f_{0}=f(r_{1},\dots,r_{n},r_{1}^{\sigma},\dots,r_{n}^{\sigma},\dots,r_{1}^{\sigma^{m}},\dots,r_{n}^{\sigma^{m}})$ and $f_{i}$ $(i\geq 1)$ are homogeneous polynomials of degree $i$ in $D\langle X_{m}\rangle$. Moreover, there exists $i_{0}$ such that $f_{i_{0}}\neq 0$. We have
$\ f(r_{1}+d_{1}t,\dots,r_{n}+d_{n}t,r_{1}^{\sigma}+d_{1}^{\sigma}t,\dots,r_{n}^{\sigma}+d_{n}^{\sigma}t,\dots,r_{1}^{\sigma^{m}}+d_{1}^{\sigma^{m}}t,\dots,r_{n}^{\sigma^{m}}+d_{n}^{\sigma^{m}}t)$
$=$ $\ \sum_{i=1}^{\infty}f_{i}(d_{1},\dots,d_{n},d_{1}^{\sigma},\dots,d_{n}^{\sigma},\dots,d_{1}^{\sigma^{m}},\dots,d_{n}^{\sigma^{m}})t^{i}.$
Since $f(r_{1}+d_{1}t,\dots,r_{n}+d_{n}t,r_{1}^{\sigma}+d_{1}^{\sigma}t,\dots,r_{n}^{\sigma}+d_{n}^{\sigma}t,\dots,r_{1}^{\sigma^{m}}+d_{1}^{\sigma^{m}}t,\dots,r_{n}^{\sigma^{m}}+d_{n}^{\sigma^{m}}t)$ is an element in the quotient ring $D(t)$ of the polynomial ring $D[t]$, it can be represented as
$f(r_{1}+d_{1}t,\dots,r_{n}+d_{n}t,r_{1}^{\sigma}+d_{1}^{\sigma}t,\dots,r_{n}^{\sigma}+d_{n}^{\sigma}t,\\ \dots,r_{1}^{\sigma^{m}}+d_{1}^{\sigma^{m}}t,\dots,r_{n}^{\sigma^{m}}+d_{n}^{\sigma^{m}}t)=\frac{g(t)}{h(t)}$
where $g(t),h(t)\in D[t]$ and $h(t)\neq 0$. Since the field $F$ is infinite, there exist infinitely many $p\in F$ such that $f$ is defined at $(r_{1}+d_{1}p,\dots,r_{n}+d_{n}p)$. Furthermore, since $f$ is a $\sigma^{m}$-GRI of $D$, we have
$f(r_{1}+d_{1}p,\dots,r_{n}+d_{n}p,r_{1}^{\sigma}+d_{1}^{\sigma}p,\dots,r_{n}^{\sigma}+d_{n}^{\sigma}p,\dots,r_{1}^{\sigma^{m}}+d_{1}^{\sigma^{m}}p,\dots,r_{n}^{\sigma^{m}}+d_{n}^{\sigma^{m}}p)=\frac{g(p)}{h(p)}=0,$
@@
$f(r_{1}+d_{1}t,\dots,r_{n}+d_{n}t,r_{1}^{\sigma}+d_{1}^{\sigma}t,\dots,r_{n}^{\sigma}+d_{n}^{\sigma}t,\dots,r_{1}^{\sigma^{m}}+d_{1}^{\sigma^{m}}t,\dots,r_{n}^{\sigma^{m}}+d_{n}^{\sigma^{m}}t)=0,$
@@
$\sum_{i=1}^{\infty}f_{i}(d_{1},\dots,d_{n},d_{1}^{\sigma},\dots,d_{n}^{\sigma},\dots,d_{1}^{\sigma^{m}},\dots,d_{n}^{\sigma^{m}})t^{i}=0.$
Therefore,
$f_{i}(d_{1},d_{2},\dots,d_{n},d_{1}^{\sigma},d_{2}^{\sigma},\dots,d_{n}^{\sigma},\dots,d_{1}^{\sigma^{m}},d_{2}^{\sigma^{m}},\dots,d_{n}^{\sigma^{m}})=0$
for all $i$. Consequently, for all $d_{1},d_{2},\dots,d_{n}\in D$, we have
$f_{i_{0}}(d_{1},d_{2},\dots,d_{n},d_{1}^{\sigma},d_{2}^{\sigma},\dots,d_{n}^{\sigma},\dots,d_{1}^{\sigma^{m}},d_{2}^{\sigma^{m}},\dots,d_{n}^{\sigma^{m}})=0.$
@@
$f_{i_{0}}(x_{1},x_{2},\dots,x_{n},x_{1}^{\sigma},x_{2}^{\sigma},\dots,x_{n}^{\sigma},\dots,x_{1}^{\sigma^{m}},x_{2}^{\sigma^{m}},\dots,x_{n}^{\sigma^{m}})$
is a $\sigma^{m}$-GPIGPI of $D$. According to Proposition 3.3, the ring $D$ satisfies a GPI. Therefore, $D$ is centrally finite by Lemma 3.1.